OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

OpenStudy (freckles):

have you tried using the formula for finding the angle between two vectors?

OpenStudy (freckles):

solve for theta \[\cos(\theta)=\frac{ u \cdot v}{|u||v|}\]

OpenStudy (anonymous):

cos(theta) = 6 * 7 / |-1| |-4| ?

OpenStudy (freckles):

how did you get that?

OpenStudy (anonymous):

i just plug in the numbers from the question to the formula you gave me. im sorry, i really dont know how to do this..

OpenStudy (freckles):

it doesn't look like you tried to find the dot product of the two vectors nor the magnitude of each vector dot product: you multiply corresponding components and then add those results: example: \[<x_1,x_2> \cdot <y_1,y_2>=x_1 y_1 +x_2 y_2\] to find the magnitude, square each component then add those results and take the square root of that sum example: \[|<x_1,x_2>|=\sqrt{x^2_1+x^2_2}\]

OpenStudy (anonymous):

dot product: 42 + 4 = 46 magnitude: 36 + 1 = 37 = 6.08 ?

OpenStudy (freckles):

\[u \cdot v=<6,-1> \cdot <7,-4>=6(7)+(-1)(-4)=42+4=46 \\ |u|=\sqrt{6^2+(-1)^2}=\sqrt{36+1}=\sqrt{37}\] almost you forgot to take the square root

OpenStudy (freckles):

you still need to find |v|

OpenStudy (anonymous):

i took the square root of 36 + 1 and got 6.08 do i find |v| then same way?

OpenStudy (freckles):

well if you mean by what I told you above yes \[|<x_1,x_2>|=\sqrt{x_1^2+x^2_2}\] this is not going to change because you have different numbers in your vector...

OpenStudy (anonymous):

49 + 16 = 65 square root = 8.06

OpenStudy (freckles):

sqrt(49+16)=sqrt(65) then this will be correct

OpenStudy (freckles):

\[u \cdot v=46 \\ |u|=\sqrt{37} \\ |v|=\sqrt{65}\]

OpenStudy (freckles):

enter this into the formula I gave above and solve for theta

OpenStudy (anonymous):

okay, so the formula is cos(theta) = u * v / |u| |v| cos(theta) - u * v / |46| |√37| what would be the values of u and v? <x1,y1> ?

OpenStudy (freckles):

you do realize the formula is: \[\cos(\theta)=\frac{ u \cdot v}{|u||v|}\] we already found u dot v in the numerator and we already found ||u| and |v| in denominator

OpenStudy (freckles):

I even summarized these values above

OpenStudy (freckles):

\[u \cdot v=46 \\ |u|=\sqrt{37} \\ |v|=\sqrt{65}\] \[\cos(\theta)=\frac{ u \cdot v}{|u||v|}\] \[\cos(\theta)=\frac{46}{\sqrt{37} \sqrt{65}}\]

OpenStudy (anonymous):

cos(theta) = 46 / √37 √65

OpenStudy (anonymous):

oh, okay.. i see

OpenStudy (freckles):

last step solve for theta theta is the angle after all

OpenStudy (anonymous):

i got 60.96..

OpenStudy (freckles):

maybe you entered something wrong I'm getting this: http://www.wolframalpha.com/input/?i=arccos%2846%2F%28sqrt%2837%29*sqrt%2865%29%29%29

OpenStudy (anonymous):

i see, so it would be 20.3 degrees

OpenStudy (freckles):

yes but you need to figure out what you are doing in your calculator so you can get that answer by yourself

OpenStudy (anonymous):

yeah i know, i found out what i did wrong.

OpenStudy (freckles):

well that's great

OpenStudy (anonymous):

thank you so much for your help! you are awesome!

OpenStudy (freckles):

it was no problem